OpenStudy (anonymous):

f(x)=3x+1 g(x)=x^2-5x-2 Find g(f(x))

OpenStudy (anonymous):

so first I plug in f(x) to replace all the x's in g(x) right?

OpenStudy (anonymous):

so I'd get g(f(x))=(3x+1)^2-5(5x+1)-2

OpenStudy (anonymous):

@Jhannybean @jim_thompson5910

OpenStudy (anonymous):

9x^2-19x-6 final answer

OpenStudy (kirbykirby):

Maybe it's easier to understand it if you re-label things. let f(x) = u Now find g(u). This meas plugging in u in g: g(u) = u^2-5u-2 But now, what is u? It's just f(x).. so replace all "u" with f(x) = 3x+1 g(u)=g(f(x))=(3x+1)^2-5(3x+1)-2

OpenStudy (anonymous):

oh whoops my bad 9x^2-9x-6

OpenStudy (kirbykirby):

^yes that's correct